# Transformers for Stratified Spectropolarimetric Inversion: Proof of Concept

### v1

@@ 6 Outlook
This study provides the first transformer-based model for full-vector Stokes inversions in solar physics, enabling accurate and efficient inference of atmospheric parameters over entire 2D fields of view and optical depths (or heights). We anticipate a follow-up study to apply the transformer to real spectropolarimetric observations, such as those from Sunrise III, DKIST, and other current or next-generation facilities. Training on simulations with advanced chromospheric physics will be critical to ensure robust generalisation. We may find that the transformer is most useful in providing a close initial guess for subsequent inversion.
R.J.C. would like to thank Robert Ryans and Ernst de Mooij for approving use of an NVIDIA Quadro RTX 8000 GPU for model development, as well as technical staff from the Kelvin2 HPC support teams in using their NVIDIA H100 GPUs. R.J.C. thanks Elena Khomenko for permission to use the MANCHA models. This is UKRI research supported research under the Science and Technology Facilities Council (STFC) grants ST/P000304/1 and ST/X000923/1. Data availability statement: The code that defined the models generated in this study is available on GitHub. The MANCHA simulation code is available on their public repository. The simulation data used in this study were generated by a third party and are not publicly archived. Access to these data may be available from the original authors on request. {contribution} R.J.C. conceived the project, designed and implemented the machine learning models, conducted the analysis and authored the paper. M.M. provided critical feedback on the framing of the study and the manuscript. C.Q.N. provided simulation data converted to an optical depth scale that formed the training and test set. All authors approved the manuscript before submission.

### v2

@@ 6 Outlook
This study provides the first transformer-based model for full-vector Stokes inversions in solar physics, enabling accurate and efficient inference of atmospheric parameters over many spectral lines and at many optical depths. We anticipate a follow-up study to apply the transformer to real spectropolarimetric observations, such as those from Sunrise III, DKIST, and other current or next-generation facilities. Training on simulations with advanced chromospheric physics will be critical to ensure robust generalisation. We may find that the transformer is most useful in providing a close initial guess for subsequent inversion.
R.J.C. would like to thank Robert Ryans and Ernst de Mooij for approving use of an NVIDIA Quadro RTX 8000 GPU for model development, as well as technical staff from the Kelvin2 HPC support teams in using their NVIDIA H100 GPUs. R.J.C. thanks Elena Khomenko for permission to use the MANCHA models. This is UKRI-supported research under the Science and Technology Facilities Council (STFC) grants ST/P000304/1 and ST/X000923/1. Data availability statement: The code that defined the models generated in this study is available on GitHub. The MANCHA simulation code is available on their public repository. The simulation data used in this study were generated by a third party and are not publicly archived. Access to these data may be available from the original authors on request. {contribution} R.J.C. conceived the project, designed and implemented the machine learning models, conducted the analysis and authored the paper. $M.$ provided critical feedback on the framing of the study and the manuscript. C.Q.N. provided simulation data converted to an optical depth scale that formed the training and test set. All authors approved the manuscript before submission.